OpenStudy (anonymous):

Find the solution set by factoring. x^2 = 10x

OpenStudy (anonymous):

x * x = 10 * x x = 10

OpenStudy (anonymous):

The solution set?

OpenStudy (anonymous):

Oy, thanks (: